# LAURA: LLM-Assisted UAV Routing for AoI Minimization

### v1

@@ Abstract
With the rapid growth of the low-altitude economy, there is increasing demand for real-time data collection using UAV-assisted wireless sensor networks. This paper investigates the problem of minimizing the age of information (AoI) in UAV-assisted wireless sensor networks by optimizing the UAV flight routing. We formulate the AoI minimization task and propose a large language model (LLM)-assisted UAV routing algorithm (LAURA). LAURA employs an LLM as intelligent crossover operators within an evolutionary optimization framework to efficiently explore the solution space. Simulation results show that LAURA outperforms benchmark methods in reducing the maximum AoI, especially in scenarios with a large number of sensor nodes.
@@ II-A Network Model
Figure 1: System model with one UAV, one data center and $N$ SNs, using LAURA which integrates an LLM and evolutionary mechanisms to optimize the UAV routing.
We consider a UAV-assisted WSN system, consisting of a data center $s_{0}$, a single UAV, and $N$ ground SNs denoted by $\bm{\mathcal{S}}=\{s_{1},s_{2},\dots,s_{N}\}$. Each SN $s_{i}$, including the date center $s_{0}$, is stationary and positioned at coordinates $(x_{i},y_{i},0)$ in a three-dimensional Cartesian space.
In such a system, the UAV serves as a mobile data aggregator, tasked with collecting data from the SNs. To facilitate a tractable analysis [10], we consider that the UAV operates at a fixed altitude $h$ and maintains a constant speed $v$. Thus, the flight time $t_{i,j}$ required for the UAV to travel between any two SNs $s_{i}$ and $s_{j}$ is given by
@@ II-B Data Transmission
When the UAV hovers above an SN $s_{i},\forall i\in\{1,2,\dots,N\}$, an uplink communication link is established, allowing the SN to transmit its data to the UAV. Since the time required for link establishment is negligible compared to the data transmission and UAV flight time [11], it is omitted from the later analysis.
Each SN transmits data to the UAV with a fixed power $P_{0}$, the uplink data transmission rate at SN $s_{i}$ is given by
$R_{i}=W\text{log}_{2}{\left(1+\tfrac{P_{0}g_{i}}{\sigma_{\textrm{r}}^{2}h^{2}}\right)},$ (3)
where $W$ is the total system bandwidth, $g_{i}$ denotes the channel gain at the reference distance [3], and $\sigma_{\textrm{r}}$ is the received noise power. Let $D_{i}$ denote the amount of data waiting to be transmitted from $s_{i}$. Thus, the data transmission time required for the UAV to fully collect data from $s_{i}$ is given by
$\tau_{i}=\tfrac{D_{i}}{R_{i}}.$ (4)
@@ II-C Age-of-Information (AoI) Metric
where $t_{i}$ is the time at which $s_{i}$ generates and uploads its data. $u(x)=\max\{x,0\}$ is the Heaviside step function ensuring that AoI is non-negative.
@@
$A_{c_{i}}=A_{c_{i}}(t_{\textrm{end}})=\sum\limits_{j={i}}^{N}\tau_{c_{j}}+\sum\limits_{j=i}^{N}t_{c_{j},c_{j+1}},$ (6)
where $\sum_{j={i}}^{N}\tau_{c_{j}}$ is the cumulative data transmission time for the SNs from $c_{i}$ onwards in the routing sequence, and $\sum_{j=i}^{N}t_{c_{j},c_{j+1}}$ represents the overall flight time required by the UAV to complete the data collection from $c_{i}$, visit all remaining SNs in the sequence, and finally return to the data center $s_{0}$.
@@ III Problem Formulation
where $t_{c_{i},c_{j}}=\frac{d_{c_{i},c_{j}}}{v}$ denotes the UAV travel time between SNs, which is based on their Euclidean distance. Constraints (7a) and (7b) represent the UAV’s data collection duration and travel time respectively, which should be greater than zero. Constraint (7c) ensures that the UAV flight starts and ends at the data center $s_{0}$, and constraint (7d) guarantees that each SN is visited only once.
@@
For any flight routing, the AoI of the SNs decreases monotonically with the data collection order $\{c_{1},c_{2},\dots,c_{N}\}$.
@@
Given that the AoI of SNs decreases monotonically with the order in which they are visited, we need to show that $A_{c_{1}}>A_{c_{2}}>\dots>A_{c_{i}}>\dots>A_{c_{j}}>\dots>A_{c_{N}},i,j\in\{1,2\dots,N\}$ with $i\neq j$ and $i<j$. Specifically, we need to prove that
@@
Considering the difference in AoI between two visits $c_{i}$ and $c_{j}$, it is obtained that
@@
Since both $\tau_{c_{k}}$ and $t_{c_{k},c_{k+1}}$ are non-negative, it follows that $A_{c_{i}}-A_{c_{j}}>0,i<j$. The proof is completed. ∎
@@ IV-1 Initialization
Initialization involves creating an initial population, which forms the basis for the subsequent evolutionary process. To avoid the need for expert knowledge, we utilize an LLM to generate this initial set. As illustrated in Fig. 2, we provide the LLM with a prompt $\Psi(\bm{\mathcal{S}})$ that contains the task description and specific initialization hints [13]. The code extracted from the LLM’s response is then executed to produce the initial population $\bm{\varPhi}$.
@@ IV-2 Evolution
After initialization, the population evolves through an iterative process aimed at refining individual quality. Over $N_{\text{g}}$ iterations, the LLM generate $N_{\text{p}}$ individuals. This evolution comprises three stages: selection, generation, and verification.
@@
Generation Stage: Using the parent set $\bm{\varTheta}_{\ell}$, we design an evolution prompt $\Psi(\bm{\mathcal{S}},\bm{\varTheta}_{\ell})$ that includes task descriptions, parent individuals, and specific hints [13]. This guides the LLM to generate offspring individuals via crossover operations, exploring the solution space efficiently. The offspring individual $\bm{\theta}_{\ell}$ is thus obtained by
@@
$\varOmega(\bm{\vartheta}_{\ell})=\sum\nolimits_{i=1}^{N}\tau_{\vartheta_{i}^{\ell}}+\sum\nolimits_{i=1}^{N}\frac{d_{\vartheta_{i}^{\ell},\vartheta_{i+1}^{\ell}}}{v},$ (14c)
@@ V Numerical Results
We provide numerical results to validate LAURA’s performance in the UAV-assisted WSN. In the simulation, the SNs are randomly and uniformly distributed within a circle area of 3000 m radius and the data center $s_{0}$ is located at the origin $\left(0,0\right)$. The UAV operates at a height of 30 m with a speed $v$ of 10 m/s. Unless otherwise stated, we set the key parameters as: UAV’s transmit power $P_{0}=0.3\,\text{W}$, data size $D_{i}=0.5\,\text{Mb}$, system bandwidth $W=1\,\text{MHz}$, channel gain at 1 meter $g_{i}=-50\text{dB}$, and noise power $\sigma_{\textrm{r}}^{2}=-110\,\text{dBm}$ [3]. For LAURA, we set a population size of $K=10$, select $N_{\textrm{p}}=5$ parent individuals, repeat maximum $M=3$ attempts and generate $N_{\text{g}}=10$ new individuals using an LLM, i.e., Deeepseek v3-671B, QwQ-32B and Deepseek R1-32B (a distilled version of Deepseek R1).
@@
Figure 2: Instance Prompts used in the initialization and evolution of LAURA: a Task description, Parent solutions, and Prompt-specific hints [13].
@@
Fig. 3(a) shows the mean of $A_{c_{1}}$ for LAURA, LEDMA, Genetic, Greedy and Random algorithms across $N=20$, $30$, and $40$ SNs, respectively. It is seen that LAURA achieves a lower maximum AoI than other algorithms. Specifically, LAURA (Deepseek-v3) reduces the mean AoI by 16.6% and 24.7% compared to the Greedy algorithm at $N=30$ and $N=40$, respectively. The performance gaps between them become large with the increment of $N$, because LAURA integrates LLM-guided operators for deep reasoning and extensive exploration, along with evolutionary mechanisms that preserve high-quality solutions through selective reproduction. Meanwhile, LEDMA (Deepseek-v3) achieves reductions of 13.4% and 19.6%, further highlighting the superior performance of LAURA in WSNs. Moreover, the Genetic algorithm underperforms due to its reliance on manual expert knowledge and even performs worse than the Greedy algorithm when $N=20$, revealing the limitations of traditional heuristic methods.
@@
To further evaluate the performance of LLMs, Fig. 3(c) compares the model hallucination rate $\bm{\epsilon}$ of LAURA and LEDMA. Specifically, LLMs in LEDMA hallucinate when the outputs fail to meet the criteria in (14a) and (14b), whereas in LAURA, hallucinations occur when the outputs do not satisfy all conditions in (14). We can see that $\bm{\epsilon}$ of LAURA is lower than that of LEDMA. Moreover, the gaps in $\bm{\epsilon}$ between LAURA (Deepseek-R1-32B) and LEDMA (Deepseek-R1-32B) are 17.8%, 22.1% and 32.8% at $N=20,30$ and 40, respectively.
@@
Fig. 4 compares the performance of the UAV routing achieved by our proposed LAURA with the other algorithms, where $N=30$. We can see that LAURA optimizes the UAV flight routing more effectively avoiding overlapping areas in Fig. 4(a). In contrast, the Genetic algorithm converges to a local optimum, resulting in a suboptimal upload sequence. LEDMA, implemented with Deepseek v3, offers slight improvements over the Genetic algorithm but does not reach the performance achieved by LAURA. The Greedy algorithm’s step-by-step nearest selection leads to suboptimal ordering, while the Random algorithm selecting SNs arbitrarily performs the worst.

### v2

@@ Abstract
With the rapid growth of the low-altitude economy, demand for real-time data collection via UAV-assisted wireless sensor networks (WSNs) is increasing. This paper studies the problem of minimizing the age of information (AoI) in UAV-assisted WSNs by optimizing the UAV flight routing. We formulate the AoI minimization task and propose a large language model (LLM)-assisted UAV routing algorithm (LAURA). LAURA employs an LLM as the intelligent crossover operators within an evolutionary optimization framework to efficiently explore the solution space. Simulation results show that LAURA outperforms benchmark methods in reducing the maximum AoI, especially in scenarios with many sensor nodes.
@@ II-A Network Model
Figure 1: System model with UAV, a data center, and $N$ SNs, using LAURA to optimize UAV routing via LLM and evolutionary mechanisms.
We consider a UAV-assisted WSN system, consisting of a data center $s_{0}$, a UAV, and $N$ ground SNs denoted by $\bm{\mathcal{S}}=\{s_{1},s_{2},\dots,s_{N}\}$. Each SN $s_{i}$, including the date center $s_{0}$, is stationary and positioned at coordinates $(x_{i},y_{i},0)$ in a three-dimensional Cartesian space. The UAV serves as a mobile data aggregator, tasked with collecting data from the SNs. For traceability, we assume that the UAV operates at a fixed altitude $h$ and maintains a constant speed $v$. The flight time $t_{i,j}$ required for the UAV to travel between any two SNs $s_{i}$ and $s_{j}$ is then given by
@@ II-B Data Transmission
Upon reaching an SN $s_{i},\forall i\in\{1,2,\dots,N\}$, the UAV establishes a communication link and begins collecting data. Since the time required for link establishment is considered negligible compared to the data transmission and UAV flight time, it is omitted from the analysis, as in [10]. We also assume a line-of-sight (LoS) dominated wireless channel between the UAV and each SN. The channel power gain is modeled as $\frac{g_{i}}{h^{2}}$, where $g_{i}$ denotes the channel gain at the reference distance. Under a fixed transmission power $P_{0}$, the achievable data rate from SN $s_{i}$ to the UAV is given by
$R_{i}=W\log_{2}\left(1+\frac{P_{0}g_{i}}{\sigma^{2}h^{2}}\right),$ (3)
where $W$ is the system bandwidth and $\sigma^{2}$ is the received noise power. Let $D_{i}$ be the size (in bits) of the data stored at $s_{i}$. Then, the duration required to fully transmit $D_{i}$ to the UAV, assuming a continuous and uninterrupted link, is
$\tau_{i}=\frac{D_{i}}{W\log_{2}\left(1+\frac{P_{0}g_{i}}{\sigma^{2}h^{2}}\right)}=\frac{D_{i}}{R_{i}},$ (4)
which characterizes the data collection latency from SN $s_{i}$ under a constant-rate and interference-free channel assumption.
@@ II-C Age-of-Information (AoI) Metric
where $t_{i}$ is the time at which $s_{i}$ generates and uploads its data. $u(x)=\max\{x,0\}$ is the Heaviside step function, ensuring that AoI is non-negative.
@@
$A_{c_{i}}=A_{c_{i}}(t_{\textrm{end}})=\sum\nolimits_{j={i}}^{N}\tau_{c_{j}}+\sum\nolimits_{j=i}^{N}t_{c_{j},c_{j+1}},$ (6)
where $\sum_{j={i}}^{N}\tau_{c_{j}}$ is the cumulative data transmission time for the SNs from $c_{i}$ onwards in the routing sequence, and $\sum_{j=i}^{N}t_{c_{j},c_{j+1}}$ represents the overall flight time required by the UAV to complete the data collection from $c_{i}$, visit all remaining SNs in the sequence, and finally return to the data center $s_{0}$. The UAV’s energy consumption is not explicitly modeled under the assumption of sufficient onboard energy or availability of recharging, and is thus omitted here.
@@ III Problem Formulation
where $t_{c_{i},c_{j}}=\frac{d_{c_{i},c_{j}}}{v}$ denotes the UAV travel time between SNs, which is based on their Euclidean distance. Constraints (7a) and (7b) respectively represent the UAV’s data collection duration and travel time, which should be greater than zero. Constraint (7c) ensures that the UAV flight starts and ends at the data center $s_{0}$, and constraint (7d) guarantees that each SN is visited only once.
@@
For any flight routing, the AoI of SNs decreases monotonically with the data collection order $\{c_{1},c_{2},\dots,c_{N}\}$.
@@
Since the AoI of each SN increases with its waiting time, we aim to show that if $i<j$, then the SN visited earlier, $c_{i}$, has a higher AoI than $c_{j}$, i.e.,
@@
To quantify it, consider the AoI difference between $c_{i}$ and $c_{j}$:
@@
where $\tau_{c_{k}}\geq 0$ and $t_{c_{k},c_{k+1}}\geq 0$ for all $k$. Hence, the summation is strictly positive for $i<j$, and we conclude that $A_{c_{i}}>A_{c_{j}},i<j$. The proof is complete.
@@ IV-1 Initialization
Initialization involves creating an initial population, which forms the basis for the subsequent evolutionary process. To avoid the need for expert knowledge, we utilize an LLM to generate this initial set. As illustrated in Fig. 2, we provide the LLM with a prompt $\Psi(\bm{\mathcal{S}})$ that contains the task description and specific initialization hints. The code extracted from the LLM’s response is then executed to produce the initial population $\bm{\varPhi}$.
@@ IV-2 Evolution
After initialization, the population evolves through an iterative process aimed at refining the quality of each individual. Over $N_{\text{g}}$ iterations, the LLM generate $N_{\text{p}}$ individuals. This evolution comprises three stages: selection, generation, and verification.
@@
Generation Stage: Using the parent set $\bm{\varTheta}_{\ell}$, we design an evolution prompt $\Psi(\bm{\mathcal{S}},\bm{\varTheta}_{\ell})$ that includes task descriptions, parent individuals, and specific hints. This guides the LLM to generate offspring individuals via crossover operations, exploring the solution space efficiently. The offspring individual $\bm{\theta}_{\ell}$ is thus obtained by
@@
$\varOmega(\bm{\vartheta}_{\ell})=\sum\nolimits_{i=1}^{N}\tau_{\vartheta_{i}^{\ell}}+\sum\nolimits_{i=1}^{N}\tfrac{d_{\vartheta_{i}^{\ell},\vartheta_{i+1}^{\ell}}}{v},$ (14c)
@@ V Numerical Results
We provide numerical results to validate LAURA’s performance in the UAV-assisted WSN. In the simulation, the SNs are randomly and uniformly distributed within a circle area of 3000 m radius and the data center $s_{0}$ is located at the origin $\left(0,0\right)$. The UAV operates at a height of 30 m with a speed $v$ of 10 m/s. Unless otherwise stated, we set the key parameters as: UAV’s transmit power $P_{0}=0.3\,\text{W}$, data size $D_{i}=0.5\,\text{Mb}$, system bandwidth $W=1\,\text{MHz}$, channel gain at 1 meter $g_{i}=-50\text{dB}$, and noise power $\sigma^{2}=-110\,\text{dBm}$ [3]. For LAURA, we set a population size of $K=10$, select $N_{\textrm{p}}=5$ parent individuals, repeat maximum $M=3$ attempts and generate $N_{\text{g}}=10$ new individuals using an LLM, i.e., Deeepseek v3-671B, QwQ-32B and Deepseek R1-32B (a distilled version of Deepseek R1).
@@
Figure 2: Instance Prompts used in the initialization and evolution of LAURA: a Task description, Parent solutions, and Prompt-specific hints.
@@
Fig. 3(a) shows the mean of $A_{c_{1}}$ for LAURA, LEDMA, Genetic, Greedy and Random algorithms across $N=20$, $30$, and $40$ SNs, respectively. It is seen that LAURA achieves a lower maximum AoI than other algorithms. Specifically, LAURA (Deepseek-v3) reduces the mean AoI by 16.6% and 24.7% compared to the Greedy algorithm at $N=30$ and $N=40$, respectively. The performance gaps between them become large with the increment of $N$, due to the integration of LLM-guided exploration and evolutionary mechanisms for preserving high-quality solutions. Meanwhile, LEDMA (Deepseek-v3) achieves reductions of 13.4% and 19.6%, demonstrating LAURA’s superiority in WSNs. Moreover, the Genetic algorithm underperforms due to reliance on manual expert knowledge and even performs worse than the Greedy one when $N=20$, revealing the limitations of traditional heuristic methods.
@@
Fig. 3(c) compares the model hallucination rate $\bm{\epsilon}$ of LAURA and LEDMA. Specifically, LLMs in LEDMA hallucinate when the outputs fail to meet the criteria in (14a) and (14b), whereas in LAURA, hallucinations occur when the outputs do not satisfy all conditions in (14). We can see that $\bm{\epsilon}$ of LAURA is lower than that of LEDMA. Moreover, the gaps in $\bm{\epsilon}$ between LAURA (Deepseek-R1-32B) and LEDMA (Deepseek-R1-32B) are 17.8%, 22.1% and 32.8% at $N=20,30$ and 40, respectively.
@@
Fig. 4 compares the UAV routing performance of the proposed LAURA algorithm against several baselines for $N=30$. We can see that LAURA optimizes the UAV flight routing more effectively, avoiding overlapping areas in Fig. 4(a). While the Genetic algorithm converges to a local optimum, yielding a suboptimal sequence. LEDMA, implemented with Deepseek v3, achieves slight improvements over the Genetic algorithm but still underperforms compared to LAURA. The Greedy approach, based on nearest selection, leads to inefficient ordering, while the Random strategy performs the worst due to its lack of coordination.